A small can of energy drink contains 9 more calories more than 4 times the number of calories in a shot of the same drink. If a small can and a shot contain a total of 154 calories, how many calories are in each drink?
Answer by macston(5194)

.
S=x=calories in shot; C=4x+9=calories in small can
.
S+C=154
x+(4x+9)=154
5x+9=154
5x-145
x=29
.
C=4x+9=4(29)+9=116+9=125
.
ANSWER A shot contains 29 calories and a small can contains 125 calories.
.
CHECK:
S+C=154
29+125=154
154=154
.